OpenStudy (anonymous):

f(x)=square root(x+4)-2

hartnn (hartnn):

@ellensmith Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) so ,\(f(x)=\sqrt{x+4}-2\) what needs to be done ?? can you please post the entire question.

OpenStudy (anonymous):

find the inverse

hartnn (hartnn):

right, so let \(y=\sqrt{x+4}-2\) interchange 'x' and 'y' \(x=\sqrt{y+4}-2\) now can you isolate 'y' from this equation ?? y=.... that will be your inverse function.

OpenStudy (anonymous):

not sure

hartnn (hartnn):

\(x=\sqrt{y+4}-2\) add 2 on both sides, what you get ?

OpenStudy (anonymous):

x+2+square root y+4

hartnn (hartnn):

you meant, \(x+2=\sqrt{y+4}\), right ? now square both sides, what you get ?

OpenStudy (anonymous):

x^2+4x+4=y+4

hartnn (hartnn):

correct! good, now subtract 4 from both sides, what you get ?

OpenStudy (anonymous):

y=x^2+4x

hartnn (hartnn):

see, how easily 'y' was isolated ... so, your inverse function is \(f^{-1}(x)=x^2+4x\) any doubts ?

OpenStudy (anonymous):

Thanks...now one more ,please solve for x 4^x=80 or 4 to the x power _80

hartnn (hartnn):

do you know logarithms ? and next time, please make a separate post for new question....

OpenStudy (anonymous):

sort a

hartnn (hartnn):

do you have options/choices for 4^x=80 in which there are log terms ??

OpenStudy (anonymous):

just have to solve for x

hartnn (hartnn):

ok, then take log on both sides. use \(\log x^n=n\log x\) so, what you get for \(\log 4^x=.... ??\)

OpenStudy (anonymous):

not sure

hartnn (hartnn):

try again.....if you notice, the exponent (n) becomes the co-efficient of log term (log x) so, what about 4^x ?

OpenStudy (anonymous):

lost

hartnn (hartnn):